9. Find the volume of each cylinder with the given dimensions. Round your answer to the nearest hundredth.
a) R=8.1 cm H= 4 cm
b) R= 9m H= 23.5m

Answer by checkley77(12844)

a) R=8.1 cm H= 4 cm
base=2pir
base=2*3.14*8.1
base=6.28*8.1
base=50.868
v=50.868*4=203.47 ans
b) R= 9m H= 23.5m
base=2pir
base=2*3.14*9
base=6.28*9
base=56.52
v=56.52*23.5=1,328.22 ans.